Five years ago, Glow Go Inc. earned $7.50 per share.  Its earnings this year were $13.20.  What was...

  1. Five years ago, Glow Go Inc. earned $7.50 per share.  Its earnings this year were $13.20.  What was the growth rate in earnings per share (EPS) over the 5-year period?
Answer #1

Ans 11.97%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
13.20 /7.50 = (1 + r%)^5
1.76 = (1 + r%)^5
r = 11.97%